OpenStudy (nottim):

How far do the dishes move during the 0.20s it takes to pull out the tablecloth? The coefficient of kinetic friction between the cloth and the dishes is μk = 0.12.

OpenStudy (anonymous):

Well we assume it has mass \(m\) I suppose. Then the normal force will be \(mg\).

OpenStudy (anonymous):

So \(\mu_kmg\) would be the force that acts on the dishes for the fifth of a second.

OpenStudy (anonymous):

This is going to be constant acceleration.

OpenStudy (nottim):

Can we assume mass to be 1?

OpenStudy (anonymous):

So we can integrate twice and get the formula: \(\Delta x =\frac 12at^2=\frac 12u_kmgt^2\)

OpenStudy (anonymous):

I suppose you could, but I believe the end result is proportional to the mass. Lighter things will move less than heavier things because the friction is different.

OpenStudy (anonymous):

wait, scratch that...

OpenStudy (anonymous):

Since \(a=u_kg\) I suppose mass is irrelevant.

OpenStudy (anonymous):

Should have said \(\Delta x =\frac 12at^2=\frac 12u_kgt^2\)

OpenStudy (anonymous):

So just plug in your numbers.

OpenStudy (nottim):

but we aren't given acceleration? i think

OpenStudy (anonymous):

gravitational acceleration... assume it is on Earth

OpenStudy (nottim):

for both a and g?

OpenStudy (anonymous):

No, \(a=u_kg\)

OpenStudy (nottim):

i see